a chemist has one solution that is 15% salkt and a second that is 25% salt how many liters of each should be mixed together to get 30 liters of a solution that is 18% salt?
Answer by mananth(16946)

percent ---------------- quantity
Base I 15 ---------------- x L
Base II 25 ------ 30 - x L
Mixture 18.00% ---------------- 30

15 x + 25 ( 30 - x ) = 18 * 30

15 x + 750 - 25 x = 540
15 x - 25 x = 540 - 750
-10 x = -210
/ -10
x = 21 L 15.00% Base I
9 L 25.00% Base II

m.ananth@hotmail.ca